how do you solve this problem:6+2(4-1)+4*9

Answer by lenny460(1073)

how do you solve this problem:6+2(4-1)+4*9

2 * 4 = 8
2 * -1 = -2
4 * 9 = 36
Therefore:
6 + 8 - 2 + 36
6 + 8 = 14
14 - 2 = 12
12 + 36 = 48

The Answer:
48



Lennox Obuong
Algebra Student
Email: obuong3@aol.com